OpenStudy (anonymous):

Solve the equation.3(2x-1)=1/2(4x-2)+2

OpenStudy (welshfella):

first distribute the 3 and the 1/2 over the parentheses

OpenStudy (welshfella):

ill do the first one for you 3(2x-1)=1/2(4x-2)+2 3*2x + 3*-2 = 1/2(4x - 2) + 2 6x - 6 = 1/2(4x - 2) + 2 now you do the same process with 1/2(4x - 2)

OpenStudy (mathmale):

Please re-type the fraction on the right side of your equation. Is your +2 part of that fraction, or is it separate from the fraction? Use parentheses so that there is no ambiguity.

OpenStudy (anonymous):

@mathmale yes it right

OpenStudy (welshfella):

also i am assume that the 1/2(4x - 1) is like